OpenStudy (anonymous):

Show an informal proof to show that the following conjecture is true Conjecture: The sum of an odd number and an even number is an odd number

OpenStudy (anonymous):

I mostly don't really know how to type a informal proof.

OpenStudy (mathstudent55):

Think of the integers. ... , -3, -2, -1, 0, 1, 2, 3, ...

OpenStudy (mathstudent55):

Some integers are even, and some are odd. If you multiply every integer by 2, are the products odd or even?

OpenStudy (anonymous):

even

OpenStudy (mathstudent55):

Correct. ... , 2(-3), 2(-2), 2(-1), 2(0), 2(1), 2(2), 2(3), ... are all even

OpenStudy (mathstudent55):

Since 2 times each integer is an even integer, we can write For every integer n, 2n is an even integer.

OpenStudy (mathstudent55):

Then by the same token, we can write For every integer n, 2n + 1 is an odd integer.

OpenStudy (anonymous):

I just really don't understand how to prove that. with an informal proof

OpenStudy (mathstudent55):

We're getting there. So far do you understand that for every integer, n, 2n is an even integer, and for every integer n, 2n + 1 is an odd integer?

OpenStudy (mathstudent55):

Now let's add the numbers. odd even What is (2n + 1) + (2n) ? (2n + 1) + (2n) = = 2n + 1 + 2n = 4n + 1 = 2(2n) + 1 Since for every integer n, 2n is an even integer, then 2(2n) must be even. Therefore, 2(2n) + 1 = 4n + 1 is odd.